CH/Agen/95/02

Agenda

Charissa Group Meeting

To be held on: Wednesday, 11 Oct 1995, 11.00 am

To be held at:  Birmingham, Orlit 2 Room

Distribution: WNC,NMC,RAC,MF,BRF,GJG,SJAH,RAH,CDJ,JK,JSL,ACM, VP,WDMR,GT,RPW,DDW,DLW,JZ, all by e-mail/server

1. Apologies
2. Minutes of 10 May 1995 (meet9501)
3. Matters Arising / Actions
4. Personnel
5. ANU beamtime submission
6. Update on EPSRC matters

Discussion on review and ANU review

7. Position on ANU beamline installation
8. Progress on MEGHA project

i Hardware (detectors, chamber, gas supply)

ii Electronics ii Readout

ii Strip detectors in MEGHA  (MF)

ii Shipment to ANU

iii Installation and commissioning

9. Planning for next ANU running period
10. Detectors stocks and new purchases
11. Computing hardware needs  (SJAH)
12. Status of spend on current grants
13. Discussion of next grants applications
14. Status of Strasbourg plans
15. Reports of other running

Orsay, Ganil, Argonne

16. WWW pages
17. Update on status of papers
18. Talks etc.
19. Time for next physics meeting
20. Any other business
21. Date of next meeting

W.N. Catford (Secretary) 9 Oct 1995